OpenStudy (anonymous):

explain how you would solve 3/4X - 7 = 8

OpenStudy (anonymous):



OpenStudy (anonymous):

same as before how can you take the 7 to the other side?

OpenStudy (anonymous):

add it to 8 3/4X - 7 + 7 = 8 + 7 3/4X = 15

OpenStudy (anonymous):

ok now?

OpenStudy (anonymous):

divide .75 by 15?

OpenStudy (anonymous):

other way

OpenStudy (anonymous):

x = 20

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

thanks